OpenStudy (anonymous):

So I worked this problem out and I just want to know if I did it correctly. I'll upload the problem and then my steps.

OpenStudy (anonymous):

This is a parallelogram ABCD, and it says AB = 2y + 1, BC = y+ 1, CD = 7x-3 and DA=3x. Find the missing values

OpenStudy (anonymous):

So here is what I did: 1. 2y + 1 = 7x - 3, y=7/2x-2 2. 7/2x-2+1=3x, x = 2 3. 2y + 1 = 7(2)-3, which got me y=5, so I got y=5 and x=2

OpenStudy (anonymous):

looks correct :)

OpenStudy (anonymous):

Yay! Thank you :)

jhonyy9 (jhonyy9):

1. why - 2 and not - 4 ?

jhonyy9 (jhonyy9):

+1 on the other side with -3 make -4 or not ?

OpenStudy (anonymous):

It just makes sense to me this way, I don't really see what you are saying.

jhonyy9 (jhonyy9):

on the first point 2y+1= 7x-3 2y = 7x-4 y = (7x-4)/2

OpenStudy (anonymous):

Ah I see, well, that's just another way to do it I guess. This is just where my brain jumps to.

jhonyy9 (jhonyy9):

so this is bc dont use where necessary parantheses

jhonyy9 (jhonyy9):

but i checked it step by step and really this is true right ,correct

jhonyy9 (jhonyy9):

so was my pleasure good luck bye bye

OpenStudy (anonymous):

Thank you very much for letting me know :)